Prof A.R. Forrester, Dept of Chemistry, University of Aberdeen, Aberdeen, Scotland AB9 1FX Prof D.H. Sleeman, Dept of Computing Science, University of Aberdeen, Aberdeen, Scotland AB9 1FX Prof A.D. Barr, Dept of Engineering, University of Aberdeen, Aberdeen, Scotland AB9 1FX

Mr J.W. Anderson, Computing Centre, University of Aberdeen, Aberdeen, Scotland AB9 1FX Prof D.J. Bounds, Dept of Computer Science and Applied Mathematics, Aston University, Aston Triangle, Birmingham, B4 7ET Prof P.E. Barker, Dept of Chemical Engineering and Applied Chemistry, Aston University, Aston Triangle, Birmingham, B4 7ET

Prof I. Bennion, Dept of Electronic Engineering and Applied Physics, Aston University, Aston Triangle, Birmingham, B4 7ET Dr P.L. Bubb, Information Systems, Aston University, Aston Triangle, Birmingham, B4 7ET Prof P.N. Bartlett, Dept of Chemistry, University of Bath, Claverton Down, Bath, BA2 7AY

Prof J.F. Eastham, Dept of Electronic and Electrical Engineering, University of Bath, Claverton Down, Bath, BA2 7AY Prof J.P. Fitch, Dept of Mathematical Sciences, University of Bath, Claverton Down, Bath, BA2 7AY Prof A. Spence, Dept of Mathematical Sciences, University of Bath, Claverton Down, Bath, BA2 7AY

Prof D.W. Bullett, Dept of Physics, University of Bath, Claverton Down, Bath, BA2 7AY Prof D.R. Boyd, School of Chemistry, Queen’s University of Belfast, Belfast, Northern Ireland, BT7 1NN Prof M. Clint, School of Electrical Engineering and Computer Science, Queen’s University of Belfast, Belfast, Northern Ireland, BT7 1NN

Prof T.N.L. Patterson, School of Mathematics and Physics, Queen’s University of Belfast, Belfast, Northern Ireland, BT7 1NN Dr W.R. McDonough, Computer Centre, Queen’s University of Belfast, Belfast, Northern Ireland, BT7 1NN Prof P.P. Edwards, School of Chemistry, University of Birmingham, Edgbaston, Birmingham, B15 2TT

Prof A. Sloman, School of Computer Science, University of Birmingham, Edgbaston, Birmingham, B15 2TT Prof J.E. Bowcock, School of Mathematics and Statistics, University of Birmingham, Edgbaston, Birmingham, B15 2TT Prof T.D. Beynon, School of Physics and Space Research, University of Birmingham, Edgbaston, Birmingham, B15 2TT

Dr J.A. Hendry, Academic Computing Service, University of Birmingham, Edgbaston, Birmingham, B15 2TT Prof J. Stephenson, Dept of Computing, University of Bradford, Richmond Road, Bradford, BD7 1DP Dr S.J. Houghton, Computing Centre, University of Bradford, Richmond Road, Bradford, BD7 1DP Prof M.N.R. Ashfold, Dept of Chemistry, University of Bristol, Tyndall Avenue, Bristol, BS8 1TH Prof D.H.D. Warren, Dept of Computer Science, University of Bristol, Tyndall Avenue, Bristol, BS8 1TH Prof D.H. Peregrine, Dept of Mathematics, University of Bristol, Tyndall Avenue, Bristol, BS8 1TH

Prof E.D.T. Atkins, Dept of Physics, University of Bristol, Tyndall Avenue, Bristol, BS8 1TH Dr L. Burbridge, Computing Service, University of Bristol, Tyndall Avenue, Bristol, BS8 1TH Prof L. Johnson, Dept of Computer Science, Brunel University, Uxbridge, UB8 3PH

Prof D.C. Imrie, Dept of Physics, Brunel University, Uxbridge, UB8 3PH Mr. A.P. Broadbent, Computer Centre, Brunel University, Uxbridge, UB8 3PH Prof C.J. Adams, Dept of Computer Science, University of Buckingham, Buckingham, MK18 1EG

Prof D. Lynden-Bell, Institute of Astronomy, University of Cambrige, Cambridge, CB2 1TN Prof A.D. Buckingham, Dept of Chemistry, University of Cambrige, Cambridge, CB2 1TN Prof R.M. Needham, Computer Laboratory, University of Cambrige, Cambridge, CB2 1TN

Dr D.F. Hartley, Computer Laboratory, University of Cambrige, Cambridge, CB2 1TN Prof G.I. Barenblatt, Department of Applied Mathematics, University of Cambrige, Cambridge, CB2 1TN Prof T. Haywood, Dept of Computing and Information Studies, University of Central England in BirminghamPerry Barr, Birmingham, B42 2SU

Prof B. Cohen, Dept of Computer Science, City University, Northampton Square, London EC1V 0HB Mr. D.R. Vinograd, Computing Sevices, City University, Northampton Square, London EC1V 0HB Dr K. Barret, Dept of Mathematical and Information Sciences, Coventry University, Priory Street, Coventry, CV1 5FB

Dr P. Dimmer, Computing Services, Coventry University, Priory Street, Coventry, CV1 5FB Prof D.E. Conway, Dept of Computer Science, De Montfort University, The Gateway, Leicester, LE1 9BH Prof J.L. Morris, Dept of Mathematics and Computer Science, University of Dundee, Dundee, DD1 4HN

Dr N.W. Carter, Computing Service, University of Dundee, Dundee, DD1 4HN Prof I.C. Colligan, Dept of Mathematics and Computer Sciences, Dundee Institute of Technology, Bell Street, Dundee, DD1 1HG Mr A.M. Morgan, Computer Centre, Dundee Institute of Technology, Bell Street, Dundee, DD1 1HG Prof R.D. Chambers, Dept of Chemistry, University of Durham, Durham, DH1 3HP Prof K.H. Bennett, Dept of Engineering and Computer Science, University of Durham, Durham, DH1 3HP Prof E.J. Squires, Dept of Mathematical Sciences, University of Durham, Durham, DH1 3HP

Prof R. Ellis, Dept of Physics, University of Durham, Durham, DH1 3HP Prof J.J. Davies, Dept of Mathematics and Physics, University of East Anglia, Norwich, NR4 7TJ Mr J.P.G. Roper, Computing Centre, University of East Anglia, Norwich, NR4 7TJ

Prof R.J. Donovan, Dept of Chemistry, University of Edinburgh, Edinburgh, EH8 9YL Prof R.M. Burstall, Dept of Computer Science, University of Edinburgh, Edinburgh, EH8 9YL Prof T.J. Lyons, Dept of Mathematics and Statistics, University of Edinburgh, Edinburgh, EH8 9YL

Prof G.S. Pawley, Dept of Physics, University of Edinburgh, Edinburgh, EH8 9YL Dr G.R. Field, Computing Services, University of Edinburgh, Edinburgh, EH8 9YL Prof K. Bowden, Dept of Chemistry and Biological Chemistry, University of Essex, Wivenhoe Park, Colchester, CO4 3SQ

Prof J.E. Doran, Dept of Computer Science, University of Essex, Wivenhoe Park, Colchester, CO4 3SQ Prof D.J. Barber, Dept of Physics, University of Essex, Wivenhoe Park, Colchester, CO4 3SQ Dr M.J. Clark, Computing Service, University of Essex, Wivenhoe Park, Colchester, CO4 3SQ

Prof E.W. Abel, Dept of Chemistry, University of Exeter, Northcote House, The Queen’s Drive, Exeter, EX4 4QJ Prof D. Partridge, Dept of Computer Science, University of Exeter, Northcote House, The Queen’s Drive, Exeter, EX4 4QJ Prof C.A. Jones, Dept of Mathematics, University of Exeter, Northcote House, The Queen’s Drive, Exeter, EX4 4QJ

Prof J.C. Inkson, Dept of Physics, University of Exeter, Northcote House, The Queen’s Drive, Exeter, EX4 4QJ Dr H.E. Davies, Computer Unit, University of Exeter, Northcote House, The Queen’s Drive, Exeter, EX4 4QJ Prog G.W. Kirby, Dept of Chemistry, University of Glasgow, Glasgow, G12 8QQ

Prof S.L. Peyton Jones, Dept of Computing Science, University of Glasgow, Glasgow, G12 8QQ Dr I.M. Barbour, Dept of Physics and Astronomy, University of Glasgow, Glasgow, G12 8QQ Prof P.G. Harper, Dept of Physics, Herriot-Watt University, Riccarton, Edinburgh EH14 4AS Mr D.R. Rundell, Computing Services, Herriot-Watt University, Riccarton, Edinburgh EH14 4AS Prof P.B. Wells, Dept of Chemistry, University of Hull, Hull, HU6 7RX Prof G.R. Brookes, Dept of Computer Science, University of Hull, Hull, HU6 7RX

Prof G. Poots, School of Mathematics, University of Hull, Hull, HU6 7RX Dr D.W. Beard, Computer Centre, University of Hull, Hull, HU6 7RX Prof D. Budgen, Dept of Computer Science, University of Keele, Staffordshire, ST5 5BG.

Prof W. Fuller, Dept of Physics, University of Keele, Staffordshire, ST5 5BG.

Dr C.L. Silk, Computer Centre, University of Keele, Staffordshire, ST5 5BG.

Prof E.F. Caldin, Dept of Chemistry, University of Kent at Canterbury, Canterbury, CT2 7NZ

Prof J.S.R. Chisholm, Dept of Mathematics and Statistics, University of Kent at Canterbury, Canterbury, CT2 7NZ Prof G. Rickayzen, Dept of Physics, University of Kent at Canterbury, Canterbury, CT2 7NZ Prof P.H. Welch, Computing Laboratory, University of Kent at Canterbury, Canterbury, CT2 7NZ

Dr F.G. Kingston, Computing Services, Kingston University, Penrhyn Road, Kingston upon Thames, KT1 2EE Prof D. Hutchinson, Dept of Computing, University of Lancaster, University House, Lancaster, LA1 4YW Prof W.M. Fairbairn, School of Physics and Materials, University of Lancaster, University House, Lancaster, LA1 4YW

Mrs Janice E. Macklin, Computing Services, University of Lancaster, University House, Lancaster, LA1 4YW Prof K.D. Bartle, School of Chemistry, University of Leeds, Leeds, LS2 9JT.

Prof P.M. Dew,

School of Computer Studies,

University of Leeds, Leeds, LS2 9JT.

Prof F.A. Goldsworthy, Dept of Applied Mathematical Studies, University of Leeds, Leeds, LS2 9JT.

Prof I.M. Ward, Dept of Physics, University of Leeds, Leeds, LS2 9JT.

Prof M. Wells,

University Computing Service, University of Leeds, Leeds, LS2 9JT.

Prof J.H. Holloway, Dept of Chemistry, University of Leicester, University Road, Leicester, LE1 7RH Prof W.A. Light, Dept of Mathematics and Computer Science, University of Leicester, University Road, Leicester, LE1 7RH Prof E.A. Davis, Dept of Physics, University of Leicester, University Road, Leicester, LE1 7RH Mr J.D. Wilson, Computer Centre, University of Leicester, University Road, Leicester, LE1 7RH Prof D.J. Schiffrin, Dept of Chemistry, University of Liverpool, Liverpool, L69 3BX Prof M.J.R. Shave, Dept of Computer Science, University of Liverpool, Liverpool, L69 3BX

Dr J.L. Schonfelder, Computing Services, University of Liverpool, Liverpool, L69 3BX Prof C. Michael, Dept of Mathematics, Applied, and Theoretical Physics, University of Liverpool, Liverpool, L69 3BX Prof K.A. McEwen, Dept of Physics, Birkbeck College, Malet Street, London, WC1E 7HX

Prof L.V.C. Rees, Dept of Chemistry, Imperial College, London, SW7 2AZ Prof R.J. Hynds, Centre for Computer Services, Imperial College, London, SW7 2AZ Prof J.T. Stuart, Dept of Mathematics, Imperial College, London, SW7 2AZ

Prof D.M. Binnie, Dept of Physics, Imperial College, London, SW7 2AZ Prof J.N. Buxton, Dept of Computing, King’s College London, Strand, London, WC2R 2LS Prof R.E. Burge, Dept of Physics, King’s College London, Strand, London, WC2R 2LS

Dr A.J. Byerley, Computer Centre, King’s College London, Strand, London, WC2R 2LS Prof M.A. MacCallum, Astronomy Unit, Queen Mary and Westfield College, Mile End Road, London, E1 4NS Prof J. Pritchard, Dept of Chemistry, Queen Mary and Westfield College, Mile End Road, London, E1 4NS

Prof Heather M. Liddell, Dept of Computer Science, Queen Mary and Westfield College, Mile End Road, London, E1 4NS Prof A.A. Carter, Dept of Physics, Queen Mary and Westfield College, Mile End Road, London, E1 4NS Prof C.J. Mitchell, Dept of Computer Science, Royal Holloway, University of London, Egham, Surrey, TW20 0EX

Prof J.H. Callomon, Dept of Chemistry, University College London, Gower Street, London, WC1E 6BT Prof S.R. Wilbur, Dept of Computer Science, University College London, Gower Street, London, WC1E 6BT Prof M.J.B. Duff, Dept of Physics, University College London, Gower Street, London, WC1E 6BT

Mr D.C. Knight, Central Computing Service, University College London, Gower Street, London, WC1E 6BT Prof F. Wilkinson, Dept of Chemistry, Loughborough University of Technology, Loughborough, LE11 3TU Prof D.J. Evans, Parallel Algorithm Research Centre, Loughborough University of Technology, Loughborough, LE11 3TU Prof C. Storey, Dept of Mathematical Sciences, Loughborough University of Technology, Loughborough, LE11 3TU Mr D.C. Hogg, Computing Services, Loughborough University of Technology, Loughborough, LE11 3TU Prof R. Grice, Dept of Chemistry, University of Manchester, Oxford Road, Manchester, M13 9PL

Prof J.R. Gurd, Dept of Computer Science, University of Manchester, Oxford Road, Manchester, M13 9PL Prof A. Donnachie, Dept of Physics, University of Manchester, Oxford Road, Manchester, M13 9PL Prof F.H. Summers, Computing Services, University of Manchester, Oxford Road, Manchester, M13 9PL

Prof B. Richards, Dept of Computation, University of Manchester Institute of Science and Technology, Sackville Street, Manchester, M60 1QD Prof B.T. Golding, Dept of Chemistry, University of Newcastle upon Tyne, Newcastle upon Tyne, NE1 7RU Prof H. Whitfield, Dept of Computing Science, University of Newcastle upon Tyne, Newcastle upon Tyne, NE1 7RU

Prof A.M. Soward, Dept of Mathematics and Statistics, University of Newcastle upon Tyne, Newcastle upon Tyne, NE1 7RU Elizabeth D. Barraclough, Computing Service, University of Newcastle upon Tyne, Newcastle upon Tyne, NE1 7RU Prof J.P. Simons, Dept of Chemistry, University of Nottingham, University Park, Nottingham, NG7 2RD

Prof C.A. Bates, Dept of Physics, University of Nottingham, University Park, Nottingham, NG7 2RD Prof P.H. Ford, Cripps Computing Centre, University of Nottingham, University Park, Nottingham, NG7 2RD Prof F. Berry, Dept of Chemistry, The Open University, Walton Hall, Milton Keynes, MK7 6AA

Dr D.J. Broadhurst, Dept of Physics, The Open University, Walton Hall, Milton Keynes, MK7 6AA Prof K.W. Morton, Computing Laboratory, University of Oxford, Wellington Square, Oxford, OX1 2JD Prof P.G.H. Sanders, Dept of Physics, University of Oxford, Wellington Square, Oxford, OX1 2JD

Mr A.R. Gay, Computing Services, University of Oxford, Wellington Square, Oxford, OX1 2JD Prof M.J. Denham, Dept of Computing, University of Plymouth, Drake Circus, Plymouth, PL4 8AA Prof H.M. Frey, Dept of Chemistry, University of Reading, Whiteknights, Reading, RG6 2AH

Annette E. Haworth, Computer Services Centre, University of Reading, Whiteknights, Reading, RG6 2AH Prof T. Evans, Dept of Physics, University of Reading, Whiteknights, Reading, RG6 2AH Prof D. Gani, Dept of Chemistry, University of St. Andrews, St. Andrews, Fife, KY16 9AJ Prof M.D. Atkinson, Dept of Mathematical and Computational Sciences, University of St. Andrews, St. Andrews, Fife, KY16 9AJ Prof E. Wyn-Jones, Dept of Chemistry and Applied Chemistry, University of Salford, Salford, M5 4WT Prof F. Walkden, Dept of Mathematics and Computer Science, University of Salford, Salford, M5 4WT

Prof C. Sozou, Dept of Applied Mathematics and Computation, University of Sheffield, Sheffield, S10 2TN Prof Gillian A. Gehring, Dept of Physics, University of Sheffield, Sheffield, S10 2TN Mr D.H. McLain, Computing Services, University of Sheffield, Sheffield, S10 2TN

Prof D.J. Tildesley, Dept of Chemistry, University of Southampton, Highfield, Southampton, SO9 5NH Prof A.J.G. Hey, Dept of Electronics and Computer Science, University of Southampton, Highfield, Southampton, SO9 5NH Dr G.J. Daniel, Dept of Physics, University of Southampton, Highfield, Southampton, SO9 5NH

Dr K.S. Heard, Computing Services, University of Southampton, Highfield, Southampton, SO9 5NH Prof H. Thimbleby, Dept of Computing Science and Mathematics, University of Stirling, Stirling, FK9 4LA Prof D.R. McGregor, Dept of Computer Science, University of Strathclyde, Glasgow, G1 1XQ

Prof J.M. Pratt, Dept of Chemistry, University of Surrey, Guildford, GU2 5XH Dr P. Waller, Computing Unit, University of Surrey, Guildford, GU2 5XH Prof S.A. Schuman, Dept of Mathematical and Computing Sciences, University of Surrey, Guildford, GU2 5XH

Prof A.R. Adams, Dept of Physics, University of Surrey, Guildford, GU2 5XH Prof H.G. Barrow, Dept of Computing Science and Artificial Intelligence, University of Sussex, Sussex House, Brighton, BN1 9RH Prof M.C.B. Hennessy, Dept of Mathematics, University of Sussex, Sussex House, Brighton, BN1 9RH

Prof A. Chalmers, Computing Service, University of Sussex, Sussex House, Brighton, BN1 9RH Prof A.A. Hashim, Dept of Applied Computing, University of Ulster, University House, Coleraine, Northern Ireland, BT52 1SA Prof M.D. Tedd, Dept of Computer Science, University College of Wales, Aberystwyth, P.O. Box 2, Aberystwyth, SY23 2AX

Prof L. Thomas, Dept of Physics, University College of Wales, Aberystwyth, P.O. Box 2, Aberystwyth, SY23 2AX Prof M.S. Baird, Dept of Chemistry, University College of North Wales, Bangor, LL57 2DG Prof J.J. O’Reilly, School of Engineering and Computer Systems, University College of North Wales, Bangor, LL57 2DG Prof R.F. Churchhouse, Dept of Computing Mathematics, University of Wales College of Cardiff, P.O. Box 920, Cardiff, CF1 3XP Prof P. Blood, Dept of Physics and Astronomy, University of Wales College of Cardiff, P.O. Box 920, Cardiff, CF1 3XP Prof J.W. Martin, Computing Centre, University of Wales College of Cardiff, P.O. Box 920, Cardiff, CF1 3XP

Prof A. Pelter, Dept of Chemistry, University College of Swansea, Singleton Park, Swansea, SA2 8PP Prof P. Townsend, Dept of Computer Science, University College of Swansea, Singleton Park, Swansea, SA2 8PP Prof I.G. Halliday, Dept of Physics, University College of Swansea, Singleton Park, Swansea, SA2 8PP

Prof M. Joseph, Dept of Computer Science, University of Warwick, Coventry, CV4 7AL Prof P.N. Butcher, Dept of Physics, University of Warwick, Coventry, CV4 7AL Mr K.S.H. Halstead, Computing Services Centre, University of Warwick, Coventry, CV4 7AL

Prof G.G. Dodson, Dept of Chemistry, University of York, York, YO1 5DD Prof G.J. Pert, Dept of Physics, University of York, York, YO1 5DD Dr K.M. Jinks, Computing Service, University of York, York, YO1 5DD